OpenStudy (anonymous):

h=12+94t-16t^2 factor the right side so i take the bggest common factor which is 2 and 2(6+47t-8t^2) and then.. ?

OpenStudy (campbell_st):

now you need to factor the quadratic here is a process for a quadratic \[ax^2 + bx + c\] multiply a and c -8 * 6 = -48 find the factors that add to 47..... 48, -1 then set up (at + factor1)(at + factor 2) ------------------------- a in your question its (-8t + 48)(-8t - 1) ------------------ -8 remove a common from the 1st binomial -8(t - 6)(-8t - 1) ----------------- - 8 cancel the common factor which leaves h = 2(t - 6)(-8t - 1) as the factored form. or you can take -1 as a common factor and write it as h = -2(t - 6)(8t + 1) which ever you choose.

OpenStudy (anonymous):

omg this is very helpful!!! i think im getting it noww!! thank you so much!!